what is the some of
6 1/6 + 2 1/3

Found 4 solutions by user_dude2008, checkley77, kmayo36, Alan3354:
Answer by user_dude2008(1862)

6 1/6 + 2 1/3 = 6 1/6 + 2 2/6 = 8 3/6 = 8 1/2

Answer by checkley77(12844)

6 1/6 + 2 1/3
37/6+7/3
(37+2*7)/6
(37+14)/6
51/6=8 1/2 or 8.5 ans.

Answer by kmayo36(1)

8 3/6= 8 1/3

Answer by Alan3354(69443)

what is the some of
6 1/6 + 2 1/3
-------------
= 6 1/6 + 2 2/6
= 8 3/6
= 8 1/2
---------
PS It's the sum, not the some.